### v1

@@ 3.3 Data Verification and Splitting
Although we require GPT-4o to generate grounded responses in the format of “<ocr> text </ocr> <bbox> x1, y1 ,x2 ,y2 </bbox>”, GPT-4o sometimes fails to follow our requirement, resulting in wrong coordinates format or missing “<bbox></bbox>”. Therefore, we implement a rule-based filter to remove these defective samples Additionally, for poster and chart data, we extract the grounded text from the generated content and compare them with the full-page parsing annotations, filtering out samples that contain incorrect grounded text. We also performed a detailed categorization of the data. As shown in Fig. 2, DOGE-Dataset includes 1.4M multi-granular parsing data and 700K diverse-formatted instruction-tuning data across three document types: poster, chart and PDF document. The multi-granular parsing data comprises four fine-grained levels (word, phrase, line, paragraph) and a full-page level. These grounded text at different granularity have precise bounding box annotation. The instruction-tuning data comprises four types: grounding, referring, grounding-and-referring, and plain Q&A. The plain Q&A is derived by removing the grounded content from a portion of the grounding data. We construct the plain Q&A subset to enhance the diversity of the data and maintain the traditional document understanding capabilities. Detailed dataset statistics are shown in Appendix D.
4 DOGE-Bench
We introduce the DOGE-Bench for the evaluation of grounding and referring capabilities of MLLMs on visual document understanding tasks.
Figure 4: In DOGE-Bench, we categorize the samples into 7 evaluation tasks, further divided into 3 types.
Task Definition. As shown in Fig. 4, we systematically construct our benchmark by categorizing our data into distinct classes based on both input and output formats. This classification helps in designing clear evaluation metrics. We divide the input formats into two categories based on the presence of bounding boxes: Grounded Question (GQ) with bounding boxes, and Plain-Text Question (PQ) without bounding boxes. The output formats are categorized into four classes:
@@ 4 DOGE-Bench
By combining two input forms and four output forms, we derive seven document referring and grounding tasks. Among these tasks, three sub-tasks primarily assess grounding capability: grounded answering for plain text questions (${G_{a}}$), grounded reasoning for plain-text questions (${G_{r}}$), and grounded open-ended answering for plain-text questions (${G_{o}}$). The plain-text answering for grounded questions (${R_{t}}$) task evaluates referring capability. The remaining tasks, grounded answering with plain-text questions (${GR_{a}}$), grounded reasoning for grounded questions (${GR_{r}}$), and grounded open-ended answering for grounded questions (${GR_{o}}$) require the integration of both grounding and referring capabilities for successful completion.
@@
Metrics. Our benchmark evaluation encompasses two aspects: grounding performance and text answer accuracy. Following the previous works in grounded captioning [65], we evaluate the grounding and text answer separately. For grounding performance, we use $F1_{all}$ score, which evaluates grounding results as a multi-label classification problem. The generated grounded text is considered correct if Intersection over Union (IoU) between its bounding box and the GT bounding box is greater than 0.5, meanwhile its text matches with the GT text. For text answer accuracy, we use exact text matching accuracy for short-answer tasks and BLEU score for long-answer tasks.
@@
Data Statistics. Our DOGE-Bench includes  2K grounding samples,  0.5K referring samples, and  15K grounding-and-referring samples. The detailed statistics is shown in Tab. 1.
Table 1: Data statistic of DOGE-Bench.
Category Grounding Grounding-and-Referring Referring Total
${G_{a}}$ ${G_{r}}$ ${G_{o}}$ ${GR_{a}}$ ${GR_{r}}$ ${GR_{o}}$ ${R_{t}}$
#Sample 700 800 636 342 427 775 464 4144
5 DOGE
@@
Overall Architecture. As illustrated in Fig. 5, our model DOGE employ a general MLLM architecture, including a vision encoder, a projector, and a large language model. In the visual encoder component, to enable the model to handle high resolution, we first search for the best aspect ratio for the input image and dynamically segment images into multiple tiles. These tiles, along with a thumbnail of the input image, are provided as input to the vision encoder. We employ pixel shuffle[7] to improve the computational efficiency of the model when processing high-resolution images. For bounding box representation, we simply discrete the continuous coordinates into discrete values from 0 to 999, avoiding the introduction of extra module or location tokens. During inference, we transfer the user-selected regions to coordinates of bounding boxes and insert them into the query for preprocessing. After obtaining the output, we utilize post-processing to overlay bounding boxes on original document images, thereby facilitating user interaction.
Figure 5: The overall architecture of DOGE.
@@
Training Strategy. We adopt a three-stage training strategy, including pre-aligning, pre-training, and fine-tuning. The pre-aligning stage focuses on aligning the feature space of vision and language rapidly. In this stage, we freeze both vision encoder and large language model, and train the projector with a relatively large learning rate. The pre-training stage aims at document parsing capabilities. We unfreeze the vision encoder and the LLM, enabling the model to recognize diverse textual content and acquire text-reading capability. In the fine-tuning stage, We train the entire model using diverse instruction-tuning data, enhancing its instruction-following ability while activating its grounding and referring capabilities during dialogue and reasoning.
Table 2: The performance evaluation of DOGE on DOGE-Bench across grounding, grounding-and-referring, and referring conducted for three image categories: Poster, Chart, and PDF Document. ${Acc}$ is for text accuracy while ${F1_{all}}$ is for grounding performance.
@@
Poster 82.5 77.5 72 65.7 43 41.6 87.5 85.8 74.5 56.6 45.6 25.8 72.5
Chart 91.5 74 70.5 43.9 32.5 46.1 79.3 60.3 65.9 57.6 39.5 54.4 61
PDF Document 80.3 71 65 45 35.9 29.1 72.6 50 67.3 58.1 37.3 14.7 19.1
Average 84.8 74.2 69.2 51.5 37.1 41.9 79.8 65.4 69.2 57.4 40.8 31.6 50.9
@@
Donut[19] <1B 67.5 11.6 61.6 30.0 18.8 54.6 41.8 43.5 74.4 93.91
DocOwl[11] 7B 62.2 38.2 42.6 30.3 26.9 60.2 57.4 52.6 111.9 188.8
UReader[61] 7B 65.4 42.2 49.5 32.8 29.4 67.6 59.3 57.6 118.4 221.7
TextMonkey[24] 9B 73.0 28.6 59.7 37.8 31.9 - 66.9 65.9 - -
Vary[56] 7B 76.3 - - - - - 66.1 - - -
TokenPacker[22] 13B 70.0 - - - - - - - - -
DocPeida[10] 7B 47.1 15.2 - - - - 46.9 60.2 - -
QwenVL[1] 9B 65.1 35.4 - - - - 65.7 63.8 - -
IXC 2.5[69] 7B 90.9 70.0 71.2 - 53.6 85.2 82.2 78.2 - 307.5
@@
InternVL2[7] 8B 91.6 74.8 - - - - 83.3 77.4 - -
DOGE 8B 91.7 70.7 70.8 40.4 58.8 84.5 83.6 76.6 145.9 332.5
Training Dataset. In the pre-aligning stage, we utilize LLaVA-558K [33] to train the projector. For the pre-training dataset, we utilize DocStruct4M [12] along with our 1.4M multi-granular parsing data to enhance basic text reading, text grounding and text referring capabilities of DOGE. For the fine-tuning data, we adopt our ground-and-refer instruction-tuning data and meticulously selected datasets to enhance model performance across a wide range of document parsing and understanding tasks, resulting in a final fine-tuning dataset of  2M samples. The detailed composition of fine-tuning dataset can be found in supplementary materials. In Appendix E, we provide a detailed description and statistics of our training data composition.
6 Experiment
Table 4: Performance comparison of multi-grained text localization on DocLocal4K dataset.
Model Text Grounding Text Recognition
word phrase line paragraph ALL word phrase line paragraph ALL
DocOwl-1.5 70.42 76.38 85.88 91.34 80.38 70.10 67.86 73.88 70.70 70.63
DOGE- 74.79 78.41 84.1 95.46 82.64 79.87 75.73 73.88 70.1 75.83
DOGE 81.85 83.58 86.88 95.67 86.64 80.00 78.85 77.60 72.72 77.88
Table 5: The impact of varying the number of tiles during training and inference. #TT denotes the maximum number of tiles during training, and #IT indicates the maximum number during inference.
#TT #IT InfoVQA PDF Document
ANLS ${Acc_{avg}}$ ${F1_{{all}_{avg}}}$
9 9 68.91 52.88 36.21
16 16 69.75 67.84 42.4
16 20 70.65 50.73 21.0
@@ 6.1 Implementation Details
DOGE utilizes the InternViT-300M-448px [6, 7] as the vision encoder and Qwen2-7B-Instruct [60] as the LLM. In the pre-aligning stage, we only train the projector and the learning rate is set to 1e-3. In the pre-training and fine-tuning stage, all model parameters are trainable. The learning rate for the vision encoder is 2e-6, while the learning rate for other components is 1e-5. Each stage is conducted for 1 epoch. In the pre-aligning stage, the batch size is set to 256, and we only use the thumbnail for vision encoder input. In the pre-training stage, the batch size is configured to 512. The number of image tiles and the max length of the input sequence to the large language model are set to 9 and 4096, respectively. In the fine-tuning stage, the batch size is adjusted to 256, and the image tile number and max input length increase to 16 and 6144. During inference, we set the image tile number to 20 for traditional understanding tasks and 16 for grounding and referring tasks.
6.2 Results on DOGE-Bench
「 Due to the scarcity of models that support document grounding and referring Q&A, we mainly present the performance of our model on DOGE-Bench, providing a performance reference for future research. We evaluate the grounding and referring capabilities of DOGE across three data types and seven tasks.
@@ 6.2 Results on DOGE-Bench
Posters have a diverse range of font types and colors, making them ideal for assessing a model’s adaptability to different font styles. DOGE shows strong performance in both grounding and referring tasks on poster data, validating its robustness in recognizing fonts with varied styles. Charts contain fine-grained and structural elements, necessitating precise text localization and understanding capabilities. Despite these challenges, DOGE still demonstrates considerable performance. PDF document data typically features high-resolution and complex content, posing challenges on both referring and grounding tasks. Although DOGE performs well in simple question answering ($G_{a}$ and $GR_{a}$) and reasoning ($G_{r}$ and $GR_{r}$), its $F1_{all}$ score in open-ended questions is relatively low. This can be attributed to the difficulty of text localization in dense text images. Some texts could not be accurately located and their bounding boxes are missed during inference, limiting the model’s ability to provide comprehensive grounded outputs. Additional quantitative results are shown in Appendix C.
@@ 6.3 Traditional Document Understanding
To assess the overall capability of DOGE, we conduct experiments on 10 traditional document understanding benchmarks, including DocVQA [39], InfographicVQA [40], DeepForm [49] and KLC [48] for document comprehension, WTQ [43] and TabFact [5] for table understanding, ChartQA [37] for chart comprehension, TextVQA [47] and TextCaps [46] for natural image interpretation, VisualMRC [50] for webpage understanding. For metrics, we use ANLS [2] for DocVQA and InfoVQA, F1 score for DeepForm and KLC, text-matching accuracy for WTQ, TabFact, TextVQA and ChartQA, CIDEr [53] for TextCaps and VisualMRC.
As detailed in Tab. 3, we observe that DOGE achieves competitive performance across all tasks. Notably, DOGE outperforms the previous state-of-the-art model IXC2.5 by +5.3% in accuracy on WTQ and +25.0% in CIDEr on VisualMRC. It is important to highlight that our primary focus is on enhancing the grounding and referencing capabilities, and we do not engage in extensive pre-training and fine-tuning on large datasets as InternVL2 [7] and IXC2.5 [69]. Despite the limited data used for training, DOGE ranks second on four datasets while remaining highly competitive, showing that DOGE maintains a strong performance across general document understanding tasks. We believe that incorporating more data for pre-training and fine-tuning could further improve DOGE’s performance on these tasks.

### v2

@@ 3.3 Data Verification and Splitting
Although we require GPT-4o to generate grounded responses in the format of “<ocr> text </ocr> <bbox> x1, y1 ,x2 ,y2 </bbox>”, GPT-4o sometimes fails to follow our requirement, resulting in wrong coordinates format or missing “<bbox></bbox>”. Therefore, we implement a rule-based filter to remove these defective samples. Additionally, for poster and chart data, we extract the grounded text from the generated content and compare them with the full-page parsing annotations, filtering out samples that contain incorrect grounded text. Furthermore, we also performed a detailed categorization of the data. As shown in Fig. 2(Left), DOGR-Dataset includes 2.1M multi-granular parsing data and 700K diverse-formatted instruction-tuning data across three document types: poster, chart, and PDF document. The multi-granular parsing data includes four fine-grained levels (word, phrase, line, paragraph) and a full-page level, with precise bounding box annotations for the grounded text at different granularities. The instruction-tuning data comprises four types: grounding, referring, grounding-and-referring, and plain Q&A, with the plain Q&A subset derived by removing the grounded content from a portion of the grounding data. This subset is created to enhance data diversity while preserving traditional document understanding capabilities. More dataset statistics and construction details are shown in Appendix D.
4 DOGR-Bench
To evaluate the grounding and referring capabilities of MLLMs on visual document understanding tasks, we introduce the DOGR-Bench in this section.
Figure 4: Left: In DOGR-Bench, we categorize the samples into 7 evaluation tasks. Right: Data statistics of DOGR-Bench.
Task Definition. As shown in Fig. 4 (Left), we systematically construct our benchmark by categorizing our data into distinct classes based on both input and output formats. This classification helps in designing clear evaluation metrics. We divide the input formats into two categories based on the presence of bounding boxes: Grounded Question (GQ) with bounding boxes, and Plain-Text Question (PQ) without bounding boxes. The output formats are categorized into four classes:
@@ 4 DOGR-Bench
By combining two input forms and four output forms, we derive 7 document grounding and referring tasks. Among these tasks, 3 tasks primarily assess grounding capability: grounded answering for plain text questions (${G_{a}}$), grounded reasoning for plain-text questions (${G_{r}}$), and grounded open-ended answering for plain-text questions (${G_{o}}$). The plain-text answering for grounded questions (${R_{t}}$) task evaluates referring capability. The remaining tasks, grounded answering with plain-text questions (${GR_{a}}$), grounded reasoning for grounded questions (${GR_{r}}$), and grounded open-ended answering for grounded questions (${GR_{o}}$) require the integration of both grounding and referring capabilities for successful completion.
@@
Metrics. Our benchmark evaluation encompasses two aspects: grounding performance and text answer accuracy. Following the previous works in grounded captioning [60], we evaluate the grounding and text answer separately. For grounding performance, we use $F1_{all}$ score, which evaluates grounding results as a multi-label classification problem. The generated grounded text is considered correct if the Intersection over Union (IoU) between its bounding box and the GT bounding box is greater than 0.5, meanwhile its text matches with the GT text. For text answer accuracy, we use exact text matching accuracy for short-answer tasks and BLEU score for long-answer tasks.
@@
Data Statistics. DOGR-Bench includes 1.8k grounding, 0.6k referring, and 1.2k grounding-and-referring samples. See detailed statistics in Fig. 4 (Right).
5 DOGR
@@
Overall Architecture. As illustrated in Fig. 5, our model DOGR employ a general MLLM architecture, including a vision encoder, a projector, and a large language model. In the visual encoder component, to enable the model to handle high resolution, we first search for the best aspect ratio for the input image and dynamically segment images into multiple tiles. These tiles, along with a thumbnail of the input image, are provided as input to the vision encoder. We employ pixel shuffle[8] to improve the computational efficiency of the model when processing high-resolution images. For bounding boxes, we simply discretize the continuous coordinates into discrete values from 0 to 999, avoiding the introduction of extra modules or location tokens. During inference, we transfer the user-selected regions to coordinates of bounding boxes and insert them into the query for preprocessing. After obtaining the output, we utilize post-processing to overlay bounding boxes on original document images, thereby facilitating user interaction.
Figure 5: The overall architecture of DOGR.
@@
Training Strategy. We adopt a three-stage training strategy, including pre-aligning, pre-training, and fine-tuning. The pre-aligning stage focuses on aligning the feature space of vision and language rapidly. We freeze both the vision encoder and the large language model, and train the projector with a relatively large learning rate. The pre-training stage aims at document parsing capabilities. We unfreeze the vision encoder and the LLM, enabling the model to recognize diverse textual content and acquire text-reading capability. In the fine-tuning stage, we train the entire model using diverse instruction-tuning data, enhancing its instruction-following ability while activating its grounding and referring capabilities.
Training Dataset. In the pre-aligning stage, we utilize LLaVA-558K [28] to train the projector. For the pre-training dataset, we utilize DocStruct4M [14] along with our 2.1M multi-granular parsing data to enhance basic text reading, text grounding and text referring capabilities of DOGR. For the fine-tuning data, we adopt our ground-and-refer instruction-tuning data and meticulously selected datasets to enhance model performance across a wide range of document parsing and understanding tasks, resulting in a final fine-tuning dataset of  2M samples. The detailed composition of fine-tuning dataset can be found in supplementary materials. In Appendix E, we provide a detailed description and statistics of our training data composition.
6 Experiment
Table 1: Performance comparison on DocLocal4K. $\dagger$ denotes the pre-trained model using parsing data with bounding boxes.
Model Text Localization Text Recognition
word phrase line paragraph ALL word phrase line paragraph ALL
Gemini 1.5 Pro[47] 6.55 3.69 4.37 5.98 5.18 4.24 2.58 3.41 16.12 11.92
Gemini 2.0 Flash 4.03 5.54 7.36 9.69 6.49 6.18 2.74 3.06 1.01 2.17
Gemini 2.5 Flash[10] 8.57 8.86 10.74 10.72 9.65 14.51 12.45 17.41 37.71 25.04
Gemini 2.5 Pro[10] 23.19 39.11 38.57 10.10 27.91 37.22 32.09 45.39 47.90 46.27
GPT-4o[16] 6.89 2.77 3.18 8.25 5.27 6.02 2.81 4.04 3.86 4.38
InternVL2.5-8B[7] 8.89 5.19 2.60 12.61 7.21 6.88 3.74 4.65 9.08 8.29
Qwen2.5-VL-7B [1] 20.84 13.84 25.25 27.01 21.51 28.68 14.82 22.5 30.51 28.98
DocOwl-1.5$\dagger$[14] 70.42 76.38 85.88 91.34 80.38 70.10 67.86 73.88 70.70 70.63
[…]
Table 2: Comparison of existing MLLMs’ and DOGR’s performance on DOGR-Bench. For clear comparison, we relax the IoU threshold to 0.1 when evaluating ${F1_{all}}$(indicated with *). In the final results for DOGR, we set the IoU threshold to 0.5.
@@ 6 Experiment
Gemini 1.5 Flash*[47] 32.7 0.7 59.0 1.1 13.6 1.7 42.8 0.5 31.8 0.3 13.5 0.1 37.5
GPT-4o mini*[16] 64.5 0.8 48.7 0.6 9.4 0.5 40.5 0.0 23.3 0.7 7.5 0.2 18.8
Gemini 1.5 Pro*[47] 77.7 9.4 62.0 5.8 13.4 6.7 46.0 5.7 29.8 2.5 17.4 3.7 37.2
Gemini 2.0 Flash* 74.2 5.2 63.5 3.1 15.2 2.9 30.0 0.5 28.5 1.7 26.2 0.6 37.5
Gemini 2.5 Flash*[10] 80.2 38.8 59.3 25.4 12.8 21.5 55.0 22.8 42.0 14.7 16.1 8.6 40.7
Gemini 2.5 Pro*[10] 61.8 30.1 47.8 17.2 13.6 14.7 44.3 17.3 36.0 10.4 19.8 8.7 36.2
GPT-4o*[16] 79.0 8.8 47.0 3.8 12.5 9.2 50.0 0.5 26.8 0.8 12.0 1.2 39.8
Qwen2-VL-7B*[53] 41.7 1.8 4.3 1.4 9.8 5.5 11.8 0.3 0.5 1.5 10.1 0.0 35.5
InternVL2-8B*[8] 51.2 2.4 26.2 0.1 10.2 1.1 26.3 0.0 12.8 0.0 14.5 0.0 17.7
InternVL2.5-8B*[7] 58.3 3.8 33.0 0.6 13.8 2.0 23.3 0.0 15.5 0.0 18.0 0.3 8.5
Qwen2.5-VL-7B*[1] 63.8 19.5 35.0 9.8 6.1 11.3 40.0 1.5 19.0 4.35 8.1 0.6 43.0
DOGR* 83.2 76.3 67.7 54.8 38.3 59.4 82.8 71.4 68.0 39.9 43.0 37.6 60.3
DOGR 83.2 73.0 67.7 52.5 38.3 57.1 82.8 66.9 68.0 38.2 43.0 34.9 60.3
     - Poster 79.5 75.8 57.0 42.6 45.6 58.9 - - - - - - 73.5
     - Chart 90.5 72.4 69.5 56.9 32.3 62.8 89.0 74.5 70.0 39.1 45.2 50.7 60.0
     - PDF Document 79.5 70.7 76.5 57.8 36.8 49.8 76.5 59.5 66.0 37.3 40.8 19.0 47.5
@@
IXC 2.5-7B[64] 7B 90.9 70.0 71.2 - 53.6 85.2 82.2 78.2 - 307.5
@@
InternVL2-8B[8] 8B 91.6 74.8 - - - - 83.3 77.4 - -
Qwen2-VL-7B[53] 8B 94.5 76.5 - - - - 83.0 84.3 - -
DOGR 8B 91.7 70.7 70.8 40.4 58.8 84.5 83.6 76.6 145.9 332.5
Table 4: Our multi-granular parsing data ${\mathrm{MG}}$ enhances the model’s basic document localization capability.
Model Text Localization Text Recognition
word phrase line paragraph word phrase line paragraph
baseline 74.79 78.41 84.1 95.46 79.87 75.73 73.88 70.10
+ ${\mathrm{MG}}$ 81.85 83.58 86.88 95.67 80.00 78.85 77.60 72.72
Table 5: Our instruction-tuning data $\mathrm{{IT}}$ improves the model’s traditional document understanding performance.
Model
Doc
VQA
Info
VQA
Deep
Form
KLC
Chart
QA
Visual
MRC
baseline 87.57 64.58 66.13 37.10 80.88 265.9
+ ${\mathrm{IT}}$ 89.24 67.45 68.89 38.32 81.92 287.25
@@ 6.1 Implementation Details
DOGR utilizes the InternViT-300M-448px [6, 8] as the vision encoder and Qwen2-7B-Instruct [58] as the LLM. In the pre-aligning stage, we only train the projector and the learning rate is set to 1e-3. In the pre-training and fine-tuning stage, all model parameters are trainable. The learning rate for the vision encoder is 2e-6, while the learning rate for other components is 1e-5. Each stage is conducted for 1 epoch. In the pre-aligning stage, the batch size is set to 256, and we only use the thumbnail for vision encoder input. In the pre-training stage, the batch size is configured to 512. The number of image tiles and the max length of the input sequence to the large language model are set to 9 and 4096, respectively. In the fine-tuning stage, the batch size is adjusted to 256, and the image tile number and max input length increase to 16 and 6144, respectively.
6.2 Doc Grounding & Referring Evaluation
Basic capability verification. We first benchmark existing MLLMs on DocLocal4K[14] to evaluate basic text localization and recognition abilities. We report IoU@0.5 for text localization and BLEU-4 score for text recognition at word, phrase, line, and paragraph levels. As quantified in Tab. 1, current SOTA MLLMs[47, 16, 64, 1] show poor performance on both tasks. This underscores the lack of basic capabilities of existing MLLMs in fine-grained document tasks, i.e., detailed text position perception and region-level text recognition. This lack of basic capabilities also limits the model’s ability to conduct further flexible grounding and referring. On the contrary, pre-training with document parsing data endows DOGR and DocOwl-1.5[13] with basic capabilities for text localization and recognition. Moreover, our high-quality and rich document parsing data further enhances the performance of DOGR.
Performance comparison on DOGR-Bench. As shown in Tab. 2, our comprehensive evaluation reveals distinct performance patterns among existing MLLMs[16, 47, 8, 7, 53, 1] on DOGR-Bench. Notably, models such as GPT-4o (79.0% Ga $Acc$) and Gemini 1.5 Pro (77.7% Ga $Acc$) demonstrate relatively strong performance in basic document understanding tasks, indicating that these models possess fundamental document understanding abilities. However, correspondingly, the lower $Acc$ of referring-related tasks reveals their limitations in accurately perceiving and understanding bounding box coordinates. More crucially, even IoU threshold for $F1_{all}$ is relaxed to 0.1, the other MLLMs’ $F1_{all}$ scores across all tasks are consistently low, which indicates deficiencies in precise document grounding. In contrast, our model achieves superior performance in both understanding and grounding capabilities. Most remarkably, its grounding capability demonstrates a substantial advancement, substantially outperforming all existing open-source and proprietary SOTA models across multiple evaluation metrics. The detailed evaluation process and results are shown in Appendix F.
@@ 6.2 Doc Grounding & Referring Evaluation
Detailed Performance of DOGR. In the last 4 rows of Tab. 2, we detail the performance of DOGR across 3 data types and 7 tasks. Posters have a diverse range of font types and colors, making them ideal for assessing a model’s adaptability to different font styles. Charts contain fine-grained and structural elements, necessitating precise text localization and understanding capabilities. PDF document data typically features high-resolution and complex content, posing challenges on both referring and grounding tasks. Despite these challenges, DOGR demonstrates strong performance, validating its effectiveness and robustness. More quantitative results are shown in Appendix C.
@@ 6.3 Traditional Document Understanding
To assess the overall capability of DOGR, we conduct experiments on 10 traditional document understanding benchmarks, including DocVQA [34], InfographicVQA [35], DeepForm [44] and KLC [43] for document comprehension, WTQ [38] and TabFact [5] for table understanding, ChartQA [32] for chart comprehension, TextVQA [42] and TextCaps [41] for natural image interpretation, VisualMRC [45] for webpage understanding. For metrics, we use ANLS [2] for DocVQA and InfoVQA, F1 score for DeepForm and KLC, text-matching accuracy for WTQ, TabFact, TextVQA and ChartQA, CIDEr [50] for TextCaps and VisualMRC.
As detailed in Tab. 3, we observe that DOGR achieves competitive performance across all tasks. Notably, DOGR outperforms the previous state-of-the-art model IXC2.5 by +5.3% in accuracy on WTQ and +25.0% in CIDEr on VisualMRC. It is important to highlight that our primary focus is on enhancing the grounding and referencing capabilities, and we do not engage in extensive pre-training and fine-tuning on large datasets as InternVL2 [8] and IXC2.5 [64]. Despite the limited data used for training, DOGR ranks second on three datasets while remaining highly competitive, showing that DOGR maintains a strong performance across general document understanding tasks. We believe that incorporating more data for pre-training and fine-tuning could further improve DOGR’s performance on these tasks.